Solve the following equation of the form 1+p^2=qz


1
Expert's answer
2021-02-25T05:03:35-0500

1+p2-qz=0

Therefore the Charpit's auxiliary equation is

"\frac{dp}{\frac{df}{dx}+p \frac{df}{dz}}=\frac{dq}{\frac{df}{dy}+q \frac{df}{dz}}=\frac{dz}{-p\frac{df}{dp}-q\frac{df}{dq}}=\frac{dx}{\frac{-df}{dp}}=\frac{dy}{-\frac{df}{dq}}"

or

"\frac{dp}{0-pq}=\frac{dq}{0-q^2}=\frac{dz}{-2p^2+qz}=\frac{dx}{-2p}=\frac{dy}{z}"


"\frac{dp}{-p}=\frac{dq}{-q}"

"p=q"

"\frac{2}{q^3}=y/z"


"q=\sqrt[3]{\frac{2z}{y}}"

"dz=pdx+qdy" "dz=\sqrt[3]{\frac{2z}{y}}dx+\sqrt[3]{\frac{2z}{y}}dy"

"z=x\sqrt[3]{\frac{2z}{y}}-\frac{-1}{3}\sqrt[3]{\frac{2z}{y^4}}"